# Unexpected Improvements to Expected Improvement for Bayesian Optimization

### v2

@@ B.1 Common initialization heuristics for multi-start gradient-descent
A variety of different heuristics have been proposed for this. The most basic one is to restart from random points uniformly sampled from the domain (for instance, scikit-optimize [Head et al., 2021] uses this strategy). However, as we have argued in this paper, acquisition functions can be (numerically) zero in large parts of the domain, and so purely random restarts can become ineffective, especially in higher dimensions and with more data points. A common strategy is therefore to either augment or bias the restart point selection to include initial conditions that are closer to “promising points”. GPyOpt [The GPyOpt authors, 2016] augments random restarts with the best points observed so far, or alternatively points generated via Thompson sampling. Spearmint [Snoek et al., 2012] initializes starting points based on Gaussian perturbations of the current best point. BoTorch [Balandat et al., 2020] selects initial points by performing Boltzmann sampling on a set of random points according to their acquisition function value; the goal of this strategy is to achieve a biased random sampling across the domain that is likely to generate more points around regions with high acquisition value, but remains asymptotically space-filling. The initialization strategy used by Trieste [Picheny et al., 2023] works similarly to the one in BoTorch, but instead of using soft-randomization via Boltzmann sampling, it simply selects the top-$k$ points. Most recently, Gramacy et al. [2022] proposed distributing initial conditions using a Delaunay triangulation of previously observed data points. This is an interesting approach that generalizes the idea of initializing “in between” observed points from the single-dimensional case. However, this approach does not scale well with the problem dimension and the number of observed data points due to the complexity of computing the triangulation (with wall time empirically found to be exponential in the dimension, see [Gramacy et al., 2022, Fig. 3] and worst-case quadratic in the number of observed points).
@@ D.1 Experimental details
All algorithms are implemented in BoTorch. The analytic EI, qEI, cEI utilize the standard BoTorch implementations. We utilize the original authors’ implementations of single objective JES [Hvarfner et al., 2022], GIBBON [Moss et al., 2021], and multi-objective JES [Tu et al., 2022], which are all available in the main BoTorch repository. All simulations are ran with 32 replicates and error bars represent ±$\pm 2$ times the standard error of the mean. We use a Matern-5/2 kernel with automatic relevance determination (ARD), i.e. separate length-scales for each input dimension, and a top-hat prior on the length-scales in $[0.01,100]$. The input spaces are normalized to the unit hyper-cube and the objective values are standardized during each optimization iteration.
@@ D.5 Multi-Objective optimization with qLogEHVI
Figure 20 compares qLogEHVI and qEHVI on 6 different test problems with 2 or 3 objectives, and ranging from 2-30 dimensions. This includes 3 real world inspired problems: cell network design for optimizing coverage and capacity [Dreifuerst et al., 2021], laser plasma acceleration optimization [Irshad et al., 2023a], and vehicle design optimization [Liao et al., 2008, Tanabe and Ishibuchi, 2020]. The results are consistent with our findings in the single-objective and constrained cases: qLogEHVI consistently outperforms qEHVI, and the gap is larger on higher dimensional problems.
Figure 20: Sequential ($q=1$) optimization performance on multi-objective problems, as measured by the hypervolume of the Pareto frontier across observed points. This plot includes JES [Tu et al., 2022]. Similar to the single-objective case, qLogEHVI significantly outperforms all baselines on all test problems.
@@ D.6 Combining LogEI with TuRBO for High-Dimensional Bayesian Optimization
In the main text, we show how LogEI performs particularly well relative to other baselines in high dimensional spaces. Here, we show how LogEI can work synergistically with trust region based methods for high-dimensional BO, such as TuRBO [Eriksson et al., 2019].
@@ D.7 Constrained Problems
  • For the tension-compression problem, LogEI found a feasible solution with value $0.0129$ after 110 evaluations compared to the $0.012681$ reported in in [Coello and Montes, 2002].
@@ D.9 Details on Multi-Objective Problems
We consider a variety of multi-objective benchmark problems. We evaluate performance on three synthetic biobjective problems Branin-Currin ($d=2$) [Belakaria et al., 2020], ZDT1 ($d=6$) [Zitzler et al., 2000], and DTLZ2 ($d=6$) [Deb et al., 2002]. As described in 5, we also evaluated performance on three real world inspired problems. For the laser plasma acceleration problem, we used the public data available at Irshad et al. [2023b] to fit an independent GP surrogate model to each objective. We only queried te surrogate at the highest fidelity to create a single fidelity benchmark.
@@ D.11 Effect of the initialization strategy
Packages and frameworks commonly utilize smart initialization heuristics to improve acquisition function optimization performance. In Figure 24, we compare simple random restart optimization, where initial points are selected uniformly at random, with BoTorch’s default initialization strategy, which evaluates the acquisition function on a large number of points selected from a scrambled Sobol sequence, and selects $n$ points at random via Boltzman sampling (e.g., sampling using probabilities computed by taking a softmax over the acquisition values  [Balandat et al., 2020]. Here we consider 1024 initial candidates. We find that the BoTorch initialization strategy improves regret for all cases, and that qLogEI, followed by UCB show less sensitivity to the choice of initializations strategy. Figure25 examines the sensitivity of qEI to the number of initial starting points when performing standard random restart optimization and jointly optimizing the $q$ points in the batch. We find that, consistent with our empirical and theoretical results in the main text, qEI often gets stuck in local minima for the Ackley test function, and additional random restarts often improve results but do not compensate for the fundamental optimality gap. The performance of qLogEI also improves as the number of starting points increases.

### v3

@@ B.1 Common initialization heuristics for multi-start gradient-descent
A variety of different heuristics have been proposed for this. The most basic one is to restart from random points uniformly sampled from the domain (for instance, scikit-optimize [35] uses this strategy). However, as we have argued in this paper, acquisition functions can be (numerically) zero in large parts of the domain, and so purely random restarts can become ineffective, especially in higher dimensions and with more data points. A common strategy is therefore to either augment or bias the restart point selection to include initial conditions that are closer to “promising points”. GPyOpt [71] augments random restarts with the best points observed so far, or alternatively points generated via Thompson sampling. Spearmint [68] initializes starting points based on Gaussian perturbations of the current best point. BoTorch [6] selects initial points by performing Boltzmann sampling on a set of random points according to their acquisition function value; the goal of this strategy is to achieve a biased random sampling across the domain that is likely to generate more points around regions with high acquisition value, but remains asymptotically space-filling. The initialization strategy used by Trieste [66] works similarly to the one in BoTorch, but instead of using soft-randomization via Boltzmann sampling, it simply selects the top-$k$ points. Most recently, Gramacy et al. [33] proposed distributing initial conditions using a Delaunay triangulation of previously observed data points. This is an interesting approach that generalizes the idea of initializing “in between” observed points from the single-dimensional case. However, this approach does not scale well with the problem dimension and the number of observed data points due to the complexity of computing the triangulation (with wall time empirically found to be exponential in the dimension, see [33, Fig. 3] and worst-case quadratic in the number of observed points).
@@ D.1 Experimental details
All algorithms are implemented in BoTorch. The analytic EI, qEI, cEI utilize the standard BoTorch implementations. We utilize the original authors’ implementations of single objective JES [38], GIBBON [63], and multi-objective JES [73], which are all available in the main BoTorch repository. All simulations are ran with 32 replicates and error bars represent ±$\pm 2$ times the standard error of the mean. We use a Matern-5/2 kernel with automatic relevance determination (ARD), i.e. separate length-scales for each input dimension, and a top-hat prior on the length-scales in $[0.01,100]$. The input spaces are normalized to the unit hyper-cube and the objective values are standardized during each optimization iteration.
@@ D.5 Multi-Objective optimization with qLogEHVI
Figure 20 compares qLogEHVI and qEHVI on 6 different test problems with 2 or 3 objectives, and ranging from 2-30 dimensions. This includes 3 real world inspired problems: cell network design for optimizing coverage and capacity [19], laser plasma acceleration optimization [40], and vehicle design optimization [56, 70]. The results are consistent with our findings in the single-objective and constrained cases: qLogEHVI consistently outperforms qEHVI, and the gap is larger on higher dimensional problems.
Figure 20: Sequential ($q=1$) optimization performance on multi-objective problems, as measured by the hypervolume of the Pareto frontier across observed points. This plot includes JES [73]. Similar to the single-objective case, qLogEHVI significantly outperforms all baselines on all test problems.
@@ D.6 Combining LogEI with TuRBO for High-Dimensional Bayesian Optimization
In the main text, we show how LogEI performs particularly well relative to other baselines in high dimensional spaces. Here, we show how LogEI can work synergistically with trust region based methods for high-dimensional BO, such as TuRBO [24].
@@ D.7 Constrained Problems
  • For the tension-compression problem, LogEI found a feasible solution with value $0.0129$ after 110 evaluations compared to the $0.012681$ reported in in [11].
@@ D.9 Details on Multi-Objective Problems
We consider a variety of multi-objective benchmark problems. We evaluate performance on three synthetic biobjective problems Branin-Currin ($d=2$) [8], ZDT1 ($d=6$) [85], and DTLZ2 ($d=6$) [17]. As described in 5, we also evaluated performance on three real world inspired problems. For the laser plasma acceleration problem, we used the public data available at Irshad et al. [41] to fit an independent GP surrogate model to each objective. We only queried te surrogate at the highest fidelity to create a single fidelity benchmark.
@@ D.11 Effect of the initialization strategy
Packages and frameworks commonly utilize smart initialization heuristics to improve acquisition function optimization performance. In Figure 24, we compare simple random restart optimization, where initial points are selected uniformly at random, with BoTorch’s default initialization strategy, which evaluates the acquisition function on a large number of points selected from a scrambled Sobol sequence, and selects $n$ points at random via Boltzman sampling (e.g., sampling using probabilities computed by taking a softmax over the acquisition values  [6]. Here we consider 1024 initial candidates. We find that the BoTorch initialization strategy improves regret for all cases, and that qLogEI, followed by UCB show less sensitivity to the choice of initializations strategy. Figure25 examines the sensitivity of qEI to the number of initial starting points when performing standard random restart optimization and jointly optimizing the $q$ points in the batch. We find that, consistent with our empirical and theoretical results in the main text, qEI often gets stuck in local minima for the Ackley test function, and additional random restarts often improve results but do not compensate for the fundamental optimality gap. The performance of qLogEI also improves as the number of starting points increases.